# Atmospheric Mass Loss from TOI-1259 A b, a Gas Giant Planet With a White Dwarf Companion

### v1

@@ Abstract
The lack of close-in Neptune-mass exoplanets evident in transit surveys has largely been attributed either to photoevaporative mass loss or high-eccentricity migration. To distinguish between these two possibilities, we investigate the origins of TOI-1259 A b, a Saturn-mass (0.4 MJ, 1.0 RJ) exoplanet lying along the upper edge of the Neptune desert. TOI-1259 A b’s close-in ( = 3.48 days) orbit and low bulk density make the planet particularly vulnerable to photoevaporation. Using transits observed in the metastable helium line, we probe the upper atmosphere of TOI-1259 A b with the Hale Telescope at Palomar Observatory and the Near-Infrared Spectrograph on Keck II. We report an excess absorption of with Palomar and a blueshifted absorption of with Keck, consistent with an extended escaping atmosphere. Fitting this signal with a Parker wind model, we determine a corresponding atmospheric mass loss rate of log() = g/s for a thermosphere temperature of K based on the Palomar absorption and log() = g/s for a thermosphere temperature of K based on the Keck absorption. This relatively low rate suggests that this planet would not have been significantly altered by mass loss even if it formed in-situ. However, the presence of a white dwarf companion, TOI-1259 B, hints that this planet may not have formed close-in, but rather migrated inward relatively late. Given the estimated parameters of the proto-white dwarf companion, we find that high-eccentricity migration is possible for the system.
@@ 1 Introduction
If HEM is common, it will have important implications for the mass loss histories of these close-in gas giants. Mass loss rates are predicted to be highest at early times, when stars are more active and planets may be inflated (from the heat of formation) and thus more vulnerable to photoevaporative mass loss (Jackson et al., 2012; Owen, 2019). We might therefore expect planets that migrated late to be able to retain more of their atmosphere than those located on close-in orbits at earlier times.
The recently-discovered planet TOI-1259 A b is an invaluable test case for distinguishing between HEM and photoevaporation scenarios. This gas giant has a mass of 0.44 MJ, a radius of 1.02 RJ, and is found on a relatively close-in ( = 3.48 days) orbit around a K dwarf (Martin et al., 2021), placing it near the upper edge of the Neptune desert (Figure 1). The planet’s low density and short orbit render it an excellent candidate for observing photoevaporation in action. The system also includes a white dwarf companion (0.56 M⊙, AU), which is estimated to have started with a mass of 1.59 M⊙ and an orbital semi-major axis of AU (assuming adiabatic mass loss) when it was on the main sequence (Martin et al., 2021; Fitzmaurice et al., 2023). The presence of the white dwarf allows for the system to have an unusually well-constrained age (4.8 , which aligns well with gyrochronology estimates), as fitting the spectral energy distribution of the white dwarf provides a robust estimate of the white dwarf cooling age. However, perhaps most enticing about the white dwarf is that when it was on the main-sequence it may have been large enough and close enough to have excited TOI-1259 A b’s orbital eccentricity (via e.g. the Kozai-Lidov mechanism) and caused it to migrate inward from a more distant formation location. The evolution of the main sequence companion into a white dwarf would have then terminated the secular effects and, combined with tidal circularization, would have led to the planet’s close-in circular orbit (Martin et al., 2021).
@@ 2.1 Palomar/WIRC
We model each transit light curve using the open-source exoplanet package (Foreman-Mackey et al., 2021a, b). Following a procedure similar to the one outlined in Vissapragada et al. (2022b), we modeled the target with a limb-darkened light curve using the open-source starry package (Luger et al., 2019). We then multiplied this model by a systematics model comprised of a linear trend in time and a linear combination of comparison star light curves, where the linear weights of the comparison stars were a free parameter with a uniform prior of that was optimized as part of the fit.
@@
The transit light curves in our fits are parameterized using the orbital period , the ratio of the planetary to stellar radius , the impact parameter , the epoch , the ratio of the semi-major axis to stellar radius , the quadratic limb darkening coefficients , , and a jitter term log() to quantify the discrepancy between the photon noise and the observed variance in the data. We use fixed limb darkening coefficients calculated using ldtk (Husser et al., 2013; Parviainen & Aigrain, 2015) and assuming a stellar effective temperature , a surface gravity and a metallicity [Fe/H] = -0.5 dex (Martin et al., 2021). We note that while we did also run our fits with free limb darkening coefficients, our retrieved parameters did not change appreciably, so we therefore adopt the fixed limb darkening coefficients for the final version of our fit. We used the pymc3 (Salvatier et al., 2016) No U-Turn Sampler (NUTS; Hoffman & Gelman, 2011) to sample the posterior probability distribution. For each of the fits to the individual nights of WIRC data, we ran four chains with 1,500 tuning steps per chain and 2,500 draws. We list the prior and posterior ranges for the full set of astrophysical model parameters for these fits in Table 1.
@@
After fitting the individual transit light curves fits for each night, we find a transit depth of for the first night and for the second night. These depths are consistent within 1, and we therefore conclude that the helium signal did not vary significantly over the two months that separated our two observations. We then jointly fit both nights of data with a common transit depth. We also simultaneously fit a light curve to the TESS data (two minute cadence, downloaded from the lightkurve package; Lightkurve Collaboration et al., 2018) from multiple sectors over a two year period to further constrain the transit shape and provide a reference transit depth that we can use to search for excess absorption in the helium bandpass. This methodology was previously used to detect an excess absorption signal in the He∗ line for HAT-P-18 b (Paragas et al., 2021), which was later confirmed by a JWST observation spanning the 1083 nm helium line (Fu et al., 2022; Fournier-Tondreau et al., 2024). We followed the same procedure as described for the single night fits, except we use free limb darkening coefficients in the TESS bandpass. In this joint fit, we ran four chains with 5,000 tuning steps per chain and 5,000 draws. We found that our joint fit with uniform limb darkening coefficients results in retrieved coefficients that are within 1 of our calculated limb darkening coefficients, and similar to the single night fits, we proceeded with the calculated, fixed limb darkening coefficients in our final model. For all fits we found that the Gelman-Rubin parameter for all sampled parameters, which indicates convergence (Gelman & Rubin, 1992). The priors and posteriors for this fit are listed in Table 1 and the corresponding transit light curves are shown in Figure 2.
@@ 2.2 Keck/NIRSPEC
After obtaining the wavelength corrected 1D spectra, we used molecfit (Smette et al., 2015) to remove telluric absorption features. We then constructed a spectral grid, in which we interpolate all spectra onto a common wavelength grid. Using a notch filter, we corrected the spectral grid for detector fringing. We then constructed a residuals spectrum by taking the log of the spectral grid and subtracting off the mean of every row and column, resulting in every pixel displaying the fractional flux variation at that wavelength and timestamp.
To track the excess absorption, we take every column (wavelength) of the residuals image and subtract the mean out-of-transit portion of the residuals, resulting in an excess absorption at every wavelength. In order to remove continuum variations, we first masked out the helium line as well as additional strong lines, and then fit a third-order polynomial to each row with respect to the column number (wavelength) before subtracting that polynomial from the residual spectrum. We then shifted the spectrum onto the planetary rest frame by computing the radial velocity of the planet relative to the star using each epoch’s barycentric Julian date. We then linearly interpolate the spectrum onto a common wavelength grid. The final excess absorption spectrum as a function of time and wavelength is shown in the top panel of Figure 3, while the bottom panel shows the average in-transit excess absorption spectrum in the planet rest frame. Figure 4 shows the band-integrated light curve in a bandpass centered on the main peak of the helium line.
@@ 3 Results & mass loss Modeling
We measure a transit depth in the helium line of with Palomar/WIRC and in the TESS light-curve fit. This corresponds to a mid-transit excess absorption of , indicating the presence of an extended escaping atmosphere on TOI-1259 A b. With Keck/NIRSPEC we measure a wavelength-resolved blueshifted excess absorption signal with a peak value of in the line core. We compare these two measurements by convolving our Keck/NIRSPEC spectrum with the Palomar/WIRC helium filter following the methodology outlined in Vissapragada et al. (2020b), and find an excess absorption of . This is consistent with the Palomar/WIRC excess absorption within approximately 1.5. However, we note that the Keck excess absorption should be biased low as it does not account for limb darkening or ingress/egress. Therefore to get a more comparable Palomar/WIRC excess absorption, we divide the WIRC light curve by the model TESS light curve to get a mid-transit relative flux ratio. We then compute the relative flux ratios at the start of ingress and egress and take the average of these values, which we then subtract from the mid-transit relative flux ratio. We obtain a modified Palomar/WIRC excess absorption of , which is within of original value and still consistent with the convolved Keck/NIRSPEC excess absorption within approximately 1.5.
@@
Using the excess absorptions derived from both our Palomar fits and Keck measurements, we place a constraint on the mass loss rate by modeling the outflow strength following the methodology in Linssen et al. (2022). We begin by defining a grid in and and generating a one-dimensional isothermal Parker wind model (Oklopčić & Hirata, 2018; Lampón et al., 2020) for each grid point. We consider mass loss rates between and g s-1, and thermosphere temperatures between K. We assume a hydrogen fraction of 0.9 (i.e., 90% hydrogen and 10% helium by number) and calculate one-dimensional density and velocity profiles for each combination of and . We then use the open-source Cloudy package (Ferland et al., 1998, 2017) for each profile to calculate the corresponding heating and cooling rates as a function of altitude and ionization structure for each . These heating and cooling rates are used to derive a new temperature structure, which we then use as the input for a new Cloudy simulation run. We repeat this process until the heating and cooling rates are balanced. We then compare the final equilibrium temperature structure to the initial profile. If these temperatures are not within 50 K of each other, we reject the model as unphysical (Linssen et al., 2022).
@@
For the remaining subset of radiatively self-consistent models, we use Cloudy to compute the density of metastable helium as a function of radial distance. Using this density profile, we then calculate the corresponding helium absorption signal during transit and integrate over the nm WIRC bandpass and the Keck NIRSPEC-Y bandpass to predict mass loss rates for the Palomar excess absorption and Keck excess absorption respectively. As shown in Figure 5, we find that our measured excess helium absorption with Palomar/WIRC is best-matched by models with a mass loss rate of g s-1 and a thermosphere temperature of K. The measured excess helium absorption with Keck/NIRSPEC is best-matched by models with a mass loss rate of g s-1 and a thermosphere temperature of K. The two mass loss rates agree within approximately 2, which is expected for excess absorption measurements that agree within 1.5.
We benchmark the mass loss and thermosphere temperature predicted from the Keck/NIRSPEC excess absorption by fitting the excess helium triplet transmission spectrum measured by Keck/NIRSPEC. Using p-winds, we fit the helium triplet by again assuming a 1D isothermal Parker wind model, with a hydrogen fraction of 0.9 and an initial mass loss rate and thermosphere temperature based on the Palomar results. We then define an atmospheric model that computes the number densities of singlet and triplet helium in the outflow as a function of distance from the planet given an initial guess of the mass loss rate, thermosphere temperature and line-of-sight wind velocity. With this information, we can then calculate the expected helium absorption signal during transit. Using lmfit (Newville et al., 2016), we employ weighted least squares minimization to determine the best fit model to the Keck/NIRSPEC helium transmission spectrum (Figure 6). We find that the best fit model predicted a line-of-sight wind velocity of -1697.4 , a 9.4 and a thermosphere temperature of K, which are consistent with the Palomar/WIRC and Keck/NIRSPEC predicted mass loss and thermosphere temperature within 1.

### v2

@@ Abstract
The lack of close-in Neptune-mass exoplanets evident in transit surveys has largely been attributed either to photoevaporative mass loss or high-eccentricity migration. To distinguish between these two possibilities, we investigate the origins of TOI-1259 A b, a Saturn-mass (0.4 MJ, 1.0 RJ) exoplanet lying along the upper edge of the Neptune desert. TOI-1259 A b’s close-in ( = 3.48 days) orbit and low bulk density make the planet particularly vulnerable to photoevaporation. We studied the upper atmosphere of TOI-1259 A b using metastable helium 1083 nm transits observed with Palomar/WIRC and Keck/NIRSPEC. We report a band-integrated excess absorption of with Palomar/WIRC and a spectroscopically-resolved blueshifted absorption of () and () with Keck/NIRSPEC. These measurements indicate the presence of an extended escaping atmosphere. Fitting these signals with a Parker wind model, we determine a corresponding atmospheric mass loss rate of log() = g/s for thermosphere temperatures between K. This relatively low rate suggests that this planet would not have been significantly altered by mass loss even if it formed in-situ. However, the presence of a white dwarf companion, TOI-1259 B, hints that this planet may not have formed close-in, but rather migrated inward relatively late. Given the estimated parameters of the proto-white dwarf companion, we find that high-eccentricity migration is possible for the system.
@@ 1 Introduction
If HEM is common, this would have important implications for the mass loss histories of these close-in gas giants. Mass loss rates are predicted to be highest at early times, when stars are more active and planets may be inflated (from the heat of formation) and thus more vulnerable to photoevaporative mass loss (Jackson et al., 2012; Owen, 2019). We might therefore expect planets that migrated late to be able to retain more of their atmosphere than those located on close-in orbits at earlier times.
The recently-discovered planet TOI-1259 A b is an invaluable test case for distinguishing between HEM and photoevaporation scenarios. This gas giant has a mass of 0.44 MJ, a radius of 1.02 RJ, and is found on a relatively close-in ( = 3.48 days) orbit around a K dwarf (Martin et al., 2021), placing it near the upper edge of the Neptune desert (Figure 1). The planet’s low density and short orbit render it an excellent candidate for observing photoevaporation in action. The system also includes a white dwarf companion (0.56 M⊙, AU), which is estimated to have started with a mass of 1.59 M⊙ and an orbital semi-major axis of AU (assuming adiabatic mass loss) when it was on the main sequence (Martin et al., 2021; Fitzmaurice et al., 2023). The presence of the white dwarf allows for the system to have an unusually well-constrained age (4.8 , which aligns well with gyrochronology estimates), as fitting the spectral energy distribution of the white dwarf provides a robust estimate of the white dwarf cooling age. However, perhaps most enticing about the white dwarf is that when it was on the main-sequence it may have been large enough and close enough to have excited TOI-1259 A b’s orbital eccentricity (via e.g. the Kozai-Lidov mechanism) and caused it to migrate inward from a more distant formation location. As the planet migrated onto a close-in orbit, the resulting tidal circularization and reduction in semi-major axis would have terminated secular effects (Martin et al., 2021).
@@ 2.1 Palomar/WIRC
We model each transit light curve using the open-source exoplanet package (Foreman-Mackey et al., 2021a, b). Following a procedure similar to the one outlined in Vissapragada et al. (2022b), we modeled the target with a limb-darkened light curve using the open-source starry package (Luger et al., 2019). We then multiplied this model by a systematics model comprised of a linear trend in time and a linear combination of comparison star light curves, where the linear weights of the comparison stars were a free parameter with a uniform prior of that was optimized as part of the fit. The wide uniform prior allows comparison stars to be assigned negative weights. This ensures that stars with weights consistent with zero are not biased towards positive values, which would artificially increase their contribution to the overall systematics model. We note that while none of the comparison stars in our model were assigned negative weights, one comparison star had a weight that was consistent with zero. Running our models again without this comparison star produces less than differences in our retrieved posteriors.
@@
The transit light curves in our fits are parameterized using the orbital period , the ratio of the planetary to stellar radius , the impact parameter , the epoch , the ratio of the semi-major axis to stellar radius , the quadratic limb darkening coefficients , , and a jitter term log() to quantify the discrepancy between the photon noise and the observed variance in the data. We use fixed limb darkening coefficients calculated using ldtk (Husser et al., 2013; Parviainen & Aigrain, 2015) and assuming a stellar effective temperature , a surface gravity and a metallicity [Fe/H] = -0.5 dex (Martin et al., 2021). We note that while we did also run our fits with free limb darkening coefficients, our retrieved parameters did not change appreciably, so we therefore adopt the fixed limb darkening coefficients for the final version of our fit. We used the pymc3 (Salvatier et al., 2016) No U-Turn Sampler (NUTS; Hoffman & Gelman, 2011) to sample the posterior probability distribution. For each of the fits to the individual nights of WIRC data, we ran four chains with 1,500 tuning steps and 2,500 draws per chain. We list the prior and posterior ranges for the full set of astrophysical model parameters for these fits in Table 1.
@@
After fitting the individual transit light curves fits for each night, we find a transit depth of for the first night and for the second night. These depths are consistent within 1, and we therefore conclude that the helium signal did not vary significantly over the two months that separated our two observations. We then jointly fit both nights of data with a common transit depth. We also simultaneously fit a light curve to the TESS data (two minute cadence, downloaded using the lightkurve package; Lightkurve Collaboration et al., 2018) from multiple sectors over a two year period to further constrain the transit shape and provide a reference transit depth that we can use to search for excess absorption in the helium bandpass. This methodology was previously used to detect an excess absorption signal in the He∗ line for HAT-P-18 b (Paragas et al., 2021), which was later confirmed by a JWST observation spanning the 1083 nm helium line (Fu et al., 2022; Fournier-Tondreau et al., 2024). We followed the same procedure as described for the single night fits, except we use free limb darkening coefficients in the TESS bandpass. In this joint fit, we ran four chains with 5,000 tuning steps per chain and 5,000 draws. We found that our joint fit with uniform limb darkening coefficients results in retrieved coefficients that are within 1 of our calculated limb darkening coefficients, and similar to the single night fits, we proceeded with the calculated, fixed limb darkening coefficients in our final model. For all fits we found that the Gelman-Rubin parameter for all sampled parameters, which indicates convergence (Gelman & Rubin, 1992). The priors and posteriors for this fit are listed in Table 1 and the corresponding transit light curves are shown in Figure 2.
@@ 2.2 Keck/NIRSPEC
After obtaining the wavelength corrected 1D spectra, we used molecfit (Smette et al., 2015) to remove telluric absorption features and obtain the spectral resolution (). We then constructed a spectral grid, in which we interpolate all spectra onto a common wavelength grid. Using a notch filter, we corrected the spectral grid for detector fringing. We then constructed a residuals spectrum by taking the log of the spectral grid and subtracting off the mean of every row and column, resulting in every pixel displaying the fractional flux variation at that wavelength and timestamp.
To track the excess absorption, we take every column (wavelength) of the residuals image and subtract the mean out-of-transit portion of the residuals, defined as the residuals preceding first contact and following fourth contact ( is considered in-transit), resulting in an excess absorption at every wavelength. Given that the residuals image is in log space, subtracting the mean out-of-transit portion of the residuals is equivalent to dividing by the geometric mean. In order to remove continuum variations, we fit a third-order polynomial to a copy of the residuals spectrum with all strong lines (including both stellar and telluric absorption lines) masked out. We fit the polynomial to each row with respect to the column number (wavelength) of the masked residuals spectrum. We then subtract the polynomial continuum fit from the original residuals spectrum. We find that in the normalized spectral timeseries there is still residual variability at the locations of telluric lines, and therefore mask them in the final residuals spectrum. We separately confirm that stellar lines, including the stellar helium line, are not variable and as a result we do not mask them in the final residuals spectrum.
As a next step, we shifted the spectrum onto the planetary rest frame by computing the radial velocity of the planet relative to the star using each epoch’s barycentric Julian date. We then linearly interpolated the spectrum onto a common wavelength grid. The final excess absorption spectrum as a function of time and wavelength is shown in the top panel of Figure 3, while the bottom panel shows the average in-transit excess absorption spectrum in the planet rest frame. Figure 4 shows the band-integrated light curve in a bandpass centered on the main peak of the helium line. We note that the pre-ingress and post-egress baseline do not deviate significantly from unity, consistent with our previous conclusion that we do not detect any statistically significant variations in the stellar helium line.
@@ 3 Results & mass loss Modeling
We measure a transit depth in the helium line of with Palomar/WIRC and in the TESS light-curve fit. This corresponds to a mid-transit excess absorption of , indicating the presence of an extended escaping atmosphere on TOI-1259 A b. With Keck/NIRSPEC we measure a wavelength-resolved excess absorption signal with a peak value of in the line core, which appears to be blueshifted. We quantify the magnitude of the blueshift by fitting a Gaussian function to the Keck excess absorption spectrum using lmfit (Newville et al., 2016) and find that the helium line is blueshifted by . In order to facilitate comparisons with the Palomar/WIRC mid-transit excess absorption, we additionally report the excess absorption signal constructed only using data between second and third contact (, see Figure 3). This excess absorption spectrum has a peak value of , and is more reflective of the behavior at mid-transit.
In the top panel of Fig. 3, there are some hints that the location of the excess absorption peak may be changing as the transit proceeds. This kind of time-evolving structure is a common feature in many 3D models (e.g., Wang & Dai, 2021; Nail et al., 2023). We evaluated the statistical significance of this apparent wavelength shift by dividing the transit in half and comparing the location of the averaged excess absorption peak during the first half ( to mid-transit) versus the second half (mid-transit to ) of the transit, as shown in Figure 5. Using weighted least squares minimization with lmfit (Newville et al., 2016), we fit a model Gaussian to determine the location of each excess absorption peak. We find that the peak of the excess absorption occurs at during the first half of the transit corresponding to a blueshift of from the main triplet helium peak, whereas the second half has a peak at , corresponding to a blueshift of from the main triplet helium peak. These two values differ by 2.8, indicating that the location of the peak excess absorption appears to shift over the course of the transit. We note that the width of these two lines ( for the first half of transit, for the second half of transit) as well as the width of the helium line for the full transit ( ) are consistent with instrumental broadening ( ). Although a more detailed 3D model analysis is beyond the scope of this paper, we provide the data behind Figures 3-5 in order to facilitate future 3D studies.
Next, we check for time variability in TOI-1259 A b’s helium signal by comparing the Keck and Palomar excess absorption measurements. We convolved our Keck/NIRSPEC spectrum with the Palomar/WIRC helium filter by first trimming the Keck spectrum around the helium feature, which we identify using the sunbather best-fit helium models (see Section 3). We then pad zeros onto either end of the trimmed spectrum out to the full wavelength range of the Palomar/WIRC filter. We then integrate the spectrum in the Palomar bandpass following the methodology outlined in Vissapragada et al. (2020b), and find an excess absorption of from the Keck excess absorption spectrum and a Palomar/WIRC-predicted excess absorption of from the Keck excess absorption spectrum. Our excess absorption, which is more reflective of the behavior at mid-transit, is consistent with the Palomar/WIRC mid-transit excess absorption of at the 1 level. We also separately create a simulated timeseries for our Keck data in the WIRC helium bandpass by convolving each individual Keck spectrum with the WIRC transmission function and adding in a white light curve based on our TESS model fits. We plot the resulting light curve in Figure 6, and compare it to our measured WIRC transit light curve. Following a similar methodology as described in Section 2.1, we fit the simulated Keck transit lightcurve using exoplanet (Foreman-Mackey et al., 2021b, a) and pymc3 (Salvatier et al., 2016) and determine a mid-transit excess absorption of . This best-fit value is nearly identical to our band-integrated excess absorption value, but has a modestly higher uncertainty. This is unsurprising, as we estimate the uncertainty in this version of the fit using the scatter in the band-integrated timeseries prior to the start of ingress modified by an error scaling term that is optimized in the fit. We find that the best-fit uncertainty is times the scatter in the pre-ingress data. For the excess absorption spectrum we set the uncertainty in each wavelength bin equal to the standard deviation of the excess absorption spectrum at wavelengths outside the helium line. We conclude that there is no significant variability in the underlying helium signal between the Keck/NIRSPEC and Palomar/WIRC measurements.
@@
Using the excess absorptions derived from both our Palomar fits and Keck measurements, we place a constraint on the mass loss rate by modeling the outflow using sunbather, following the methodology in Linssen et al. (2022, 2024). We begin by defining a grid in and and generating a one-dimensional isothermal Parker wind model (Oklopčić & Hirata, 2018; Lampón et al., 2020; Dos Santos et al., 2022) for each grid point. We consider mass loss rates between and g s-1, and thermosphere temperatures between K. We assume a hydrogen fraction of 0.9 (i.e., 90% hydrogen and 10% helium by number) and calculate one-dimensional density and velocity profiles for each combination of and . We then use the open-source Cloudy package (Ferland et al., 1998, 2017) for each profile to calculate the corresponding heating and cooling rates as a function of altitude and ionization structure for each . These heating and cooling rates are used to derive a new temperature structure, which we then use as the input for a new Cloudy simulation run. We repeat this process until the heating and cooling rates are balanced. We then calculate the mean and standard deviation of the temperature profile in the helium line-forming region. We reject models with a large difference between this mean temperature and as unphysical (Linssen et al., 2022).
@@
For the remaining subset of radiatively self-consistent models, we use Cloudy to compute the density of metastable helium as a function of radial distance. Using this density profile, we then calculate the corresponding helium absorption signal during transit and integrate over the nm WIRC bandpass and separately convolve with the resolution of Keck NIRSPEC-Y at the metastable helium wavelengths to predict mass loss rates for the Palomar excess absorption and Keck excess absorption respectively. As shown in Figure 7, we find that our measured excess helium absorption with Palomar/WIRC is best-matched by models with a mass loss rate of g s-1 and a thermosphere temperature of K. The measured excess helium absorption with Keck/NIRSPEC is best-matched by models with a mass loss rate of g s-1 and a thermosphere temperature of K, while the measured excess helium absorption with Keck/NIRSPEC is best-matched by models with a mass loss rate of g s-1 and a thermosphere temperature of K. Both Keck/NIRSPEC mass loss rates are within approximately 1 of the Palomar mass loss rate. We note that the Keck/NIRSPEC mass loss rate is closer to the measured Palomar/WIRC mass loss rate than the value. This is likely because the average excess absorption calculated when the planet is fully in front of the star () is a closer match to the value obtained by fitting the full light curve with a transit model.
Given the unknown stellar spectrum of the host star, these mass loss rates are estimated using the proxy MUSCLES (Youngblood et al., 2017) stellar spectrum of HD 40307, which proves a close match in effective temperature, surface gravity, spectral type, metallicity and age to TOI-1259 A. However, an optical spectrum of TOI-1259 A obtained with the High Resolution Infrared Spectrograph (HIRES) on Keck (see Figure 8) reveals a log of -4.5 (Howard Isaacson, private communication), indicating an active chromosphere. TOI-1259 A may be quite active for its age, and falls near the edge of the rotation-log for K stars (Toledo-Padrón et al., 2019). Given the apparent enhanced stellar activity of the star, we repeat our mass loss models using the more active proxy stellar spectrum of Eridani (Youngblood et al., 2017). We find a maximum mass loss rate of g s-1 and a thermosphere temperature of K. We note that the true mass loss rate is likely in between our mass loss rates derived using HD 40307 and Eridani, as the log measurement is based on a single epoch of HIRES observation and if the star has a strong activity cycle or rotational modulation, the HIRES observation could have occurred near a maximum of either scenario, driving the activity of the star to appear higher than it would with repeat observations.